OpenStudy (hollz32):

Wesley draws a card at random from a well-shuffled deck of playing cards.The probability that Wesley draws a number card or a black card is.

OpenStudy (hollz32):

ik nothing of cards

OpenStudy (welshfella):

there are 40 number cards out of a pack of 52 cards. The number of black cards is 26.

OpenStudy (hollz32):

would i add the two together to get the out of numbered or black cards?

OpenStudy (welshfella):

you have to add probabilities but some black cards are also numbered . I'll have to check this.

OpenStudy (hollz32):

okay

OpenStudy (welshfella):

Yes we need to use the following rule required probability = P(Black) + P(number) - p ( black intersect number) the last part means the probability of them both occurring I couldn't find the symbol for intersect on equation editor. It looks like an inverted U.

OpenStudy (welshfella):

P(Number) = 40/52 = 10/13 P(Black) = 26/52 = 1/2

OpenStudy (welshfella):

What is the Probability of getting a card which is both numbered and black?

OpenStudy (hollz32):

i got 33/26 0r 1 7/26

OpenStudy (welshfella):

How many are numbered and black?

OpenStudy (welshfella):

how did you get that?

OpenStudy (welshfella):

P(black) + P(number) = 33/26 but you have to subtract the 'intersect' propbability

OpenStudy (welshfella):

there are 20 cards which are black and numbers so what is the probability of picking one of theses?

OpenStudy (welshfella):

20/52 agreed?

OpenStudy (welshfella):

or 10/26 so we have 20/26 + 13/26 - 10/26

OpenStudy (welshfella):

follow?

OpenStudy (welshfella):

I cant believe the equation editor hasnt got a symbol for intersecting of sets.

OpenStudy (welshfella):

unless I've missed it...

OpenStudy (hollz32):

i didnt know to do all that. so it was wrong?